A 22cm -diameter circular saw blade has a mass of 0.87kg , distributed uniformly as in a disk

Part A

What is its rotational kinetic energy at 3500rpm ?

Express your answer using two significant figures.

Part B

What average power must be applied to bring the blade from rest to 3500rpm in 3.5s ?

Express your answer using two significant figures.
Answer #1

radius of the disk, r = d/2 = 22/2 = 11 cm = 0.11 m

moment of inertia of disk, I = 0.5*M*r^2

= 0.5*0.87*0.11^2

= 5.26*10^-3 kg.m^2

w = 3500 rpm

= 3500*2*pi/60

= 366.33 rad/s

A)
Rotational kientic enrgy = 0.5*I*w^2

= 0.5*5.26*10^-3*366.33^2

= 352.94 J <<<<<<<-----Answer

B)

Average power = Energy supplied/time

= 353.94/3.5

= 101.13 Watts <<<<<<<-----Answer